Lindiwe decides that she would like to buy her daughter mbali a car when she turns 21years in 6 year's time. she deposit 6000 each month into an account earnings 8.94% interest per year compounded monthly.The amount that lindiwe (rounded to the nearest rand) will have available 6 years from now is

1. 432000

2.335.896

3. 573187

4. 568948


1
Expert's answer
2022-03-01T12:28:24-0500

Future value of ordinary annuity formula is used

"FV=A(\frac{(1+r)^n-1}{r})"

Where r=8.94/12=0.745%; A=6,000; and n=72;

"FV=6000(\frac{(1+0.00745)^{72}-1}{0.00745})=568,948"

At end of six years, Lindiwe will have 568,948 rand's.